OpenStudy (paperbacon):

@ganeshie8 please help. Will fan and medal!!!

OpenStudy (paperbacon):

use the unit circle to find the inverse function value in degrees tan -1 sqrt 3

OpenStudy (anonymous):

I can help you

OpenStudy (paperbacon):

120 90 60 30

OpenStudy (paperbacon):

Okay @jameel.saeed

ganeshie8 (ganeshie8):

\[\tan\theta = \dfrac{\sin \theta}{\cos\theta}\] notice that \(\large \dfrac{~~\frac{\sqrt{3}}{2}~~}{\frac{1}{2}} = \sqrt{3}\)